OpenStudy (anonymous):

In a triangle, the measure of angle A is 3times the measure of angle B, and the measure of angle C is 5 degrees more than the measure of angle B. What is the measure of angle A? answers: 60 degrees 75degrees 90 degrees 105 degrees... how is this problem done and what is the answer? Thank you

OpenStudy (anonymous):

i think this question is missing out a key information. Does A and B and C reside somewhere, maybe in a triangle or something?

OpenStudy (anonymous):

yes, a triangle.

OpenStudy (anonymous):

I'm going to go out on a limb here and assume we're working with a triangle. In all triangles, the angles add up to 180. That means we have a few equations we can work out:\[A+B+C=180\] as well as \[C=B+5\] and \[A = 3B\] Hope that helps!

OpenStudy (anonymous):

ah, sorta, still have to plug numbers in i dont know where though

OpenStudy (anonymous):

A triangle is equal to 180, 180 minus angle c (5) is 175, and angle b is 3x so divide 175 by 3, you get 58.3, so i'm guessing it's 60 degrees cause you can end up rounding the 58.3 to that by turning it into a fraction. I'm pretty sure that's how you do this I just finished this class.

OpenStudy (anonymous):

the answer says 105.

OpenStudy (anonymous):

The I have no idea how I got an A in that class, lol sorry.

OpenStudy (anonymous):

good for you, it's okay.

OpenStudy (anonymous):

set B=x A 3 times the size of be =3x c is 5 more than b= x+5 3x+x +(x+5)=180 just solve for x and substitute into 3x, x, and x+5

OpenStudy (anonymous):

okay, ILl try